### v2

@@ 3 Gamma-ray detection for gamma-ray spectroscopy
At the RIBF, secondary beams with incident energies of 100 to 250 MeV$/u$ are used to populate excited states of nuclei far from stability. While such energies allow the use of thicker targets, and therefore increase the luminosity, they also present different experimental challenges. Of particular importance are the large Doppler shift of the detected $\gamma$-rays, and the background generated by either atomic or nuclear processes.
@@
where $\beta$ denotes the velocity of the beam and $\theta_{\gamma}$ is the angle between the particle direction and the $\gamma$-ray emission in the laboratory frame of reference. The energy resolution that can be achieved is therefore limited by the uncertainties in the velocity, $\Delta\beta$, the emission angle of the $\gamma$-ray, $\Delta\theta_{\gamma}$, and the intrinsic resolution of the detection material, $\Delta E_{\gamma}$. In general, the use of thicker secondary targets leads to a higher energy loss and therefore a larger contribution to $\Delta\beta$. The use of MINOS Obertelli_EPJA_2014 overcomes this problem by tracking the protons ejected from the target after knockout reactions and reconstructing the interaction position. In this way thicker targets can be used without affecting the achievable energy resolution as shown in Fig. 8. While the intrinsic resolution, $\Delta E_{\gamma}$, is a property of the detector material, the angular resolution, $\Delta\theta_{\gamma}$, can be improved either by reducing the size of the detectors or by implementing a position-sensitive readout with a sub-centimeter resolution for the first interaction of the $\gamma$-rays. Figure 8 shows the effect of using the HYPATIA array (described below), which has a better intrinsic resolution than DALI2 takeuchi2014dali2 ; Murray_RAPR_2018 , as well as smaller detectors.
Figure 8: Total energy resolution for different $\gamma$-ray spectrometers. For a thick target, a large $\Delta\beta$ is obtained. This uncertainty is reduced by using MINOS. The use of smaller detectors with better intrinsic resolution, as the HYPATIA array, further improve the achievable energy resolution. The calculations were performed for $\beta=0.6$ and a 1 MeV $\gamma$-ray energy assuming a square-root dependence of the intrinsic energy resolution
@@ 3.1 Towards a new scintillation detector array: The HYPATIA project
With this requirement in mind, the HYPATIA (HYbrid Photon detector Array To Investigate Atomic nuclei) array has been designed hypatia_web . This array, depicted in Fig. 9, consists of 624 Cerium Bromide (CeBr3) detectors of $30\times 30\times 80$ mm2 forming a barrel around the target and 384 Cerium-doped Gadolinium Aluminum Gallium Garnet (HR-GAGG) detectors of $25\times 25\times 100$ mm2 placed as a forward wall. The CeBr3 detectors were chosen due to their improved energy- and time-resolution compared to NaI(Tl), their reasonable cost, and the lack of intrinsic radiation. On the other hand, the HR-GAGG crystals are non-hygroscopic and have a high effective atomic number, therefore offering the possibility to minimize the dead space and increase the efficiency and peak-to-total ratio at forward angles. Thanks to the cuboidal shape of all the detectors, it is possible to arrange them in different configurations, including surrounding the MINOS or STRASSE targets, as well as to place the array at the different spectrometers of the RIBF. Furthermore, the selected shapes and sizes of the detectors are compatible with the mechanical structure of the DALI2+ array, therefore it will be possible to gradually replace the detectors without interrupting the physics program.
@@
Figure 10: Expected Doppler-corrected $\gamma$ spectrum of 60Ca for a 150-hours measurement with HYPATIA.
The Ca isotopic chain is particularly interesting for nuclear structure studies. While the proton shell closure at $Z=20$ seems to be maintained, new neutron shell closures at $N=32$ and $N=34$ have been evidenced by spectroscopy Huck_PhysRevC.31.2226 ; Steppenbeck_Nature_2013 , mass measurements Wienholtz_Nature2013 ; Michimasa_PhysRevLett.121.022506 , and knockout reaction cross sections Gade_PhysRevC.74.021302 ; Chen_PhysRevLett.123.142501 . The possibility of a further magic number at the $N=40$ harmonic-oscillator shell closure has been widely studied Gade_physics3040077 ; Brown_physics4020035 . Although measurements on 58Ca Chen_PLB_2023138025 and 62Ti Cortes_PLB_2020135071 , the closest neighbours of 60Ca, show no indication of a shell closure at $N=40$, only the first spectroscopy of 60Ca can provide definitive answers. This measurement will be possible with the HYPATIA array, as shown in Fig. 10, where the simulated $\gamma$-ray spectrum obtained in a 150 h measurement is shown.
@@ 3.2 Future projects with Ge detectors
The $\gamma$-ray tracking technique of Ge detectors is the state-of-the-art method to achieve the best energy resolution for $\gamma$-rays. This technique has been implemented in the latest large-scale Ge arrays GRETINA PASCHALIS201344 and AGATA AKKOYUN201226 . During 2020 and 2021, the RIBF hosted the HiCARI array Wimmer2021 to conduct the first studies with tracking-type germanium detectors, which yielded successful outcomes. HiCARI was organized as a one-time campaign of experiments at the RIBF by gathering Ge detectors from other facilities. It is desirable to make continued efforts to push in-beam $\gamma$-ray spectroscopy based on tracking technology forward. For this reason, the Gamma-ray Tracking in R5 (GT-5) project was launched in 2023 at the RIKEN Nishina Center as a common and long-term platform for tracking-type Ge detectors at the RIBF to drive physics experiments and instrumentation for tracking technologies. The GT-5 has its own scientific and development programs related to in-beam $\gamma$-ray spectroscopy using RIBs. On a hardware basis, the GT-5 presently encompasses two in-house tracking detectors. One is the Compton camera telescope GREI motomura_IEEE_2007 , and the other is a position-sensitive strip germanium telescope SGT Suzuki_RAPR , both based on planar crystals with stripped segments. Furthermore, a GRETINA quad-type module was procured in 2023. In addition to the detectors, a maintenance facility for Ge detectors, electronics, and infrastructure is arranged. A unique aspect of the GT-5 is that it is not only involved in the scientific activities of the RIBF but also the framework of the TRIP (Transformative Research Innovation Platform of RIKEN platforms) project, a RIKEN-wide initiative starting in 2023. One of the main goals of the TRIP use case is to establish an optical model for radioactive isotopes that are needed for reaction analyses and predictions but have yet to be studied based on systematic experimental data. To provide systematic data to optimize optical model potential parameters, the GT-5 is teamed with the elastic scattering measurement program MESA (Measurement of Elastic Scattering Anytime anywhere on any beam) and the reaction cross-section measurement program S3CAN (Symbiotic Systematic and Simultaneous Cross-section measurements for All over the Nuclear chart). As the first opportunities for data collection, experiments with stable beams at the RIPS and with RIBs of the BigRIPS will be conducted in 2024.
@@ 4 Neutron detection for invariant-mass spectroscopy
The upgrade plans of the RIBF will enable to reach of further neutron-rich isotopes, which are expected to have various weakly bound multi-neutron systems. To extend the sensitivity of multi-neutron reconstructions, a project to extend the detector volume for better detection efficiency, NEBULA Plus, is in progress through a collaborative effort between Japan and LPC Caen in France, as part of the French-funded EXPAND project expand . The NEBULA Plus array consisting of 90 new modules has been integrated with the existing array, effectively doubling the number of detection layers. This expansion increases the one-neutron detection efficiency by nearly a factor of 2 and improve the four-neutron detection efficiency by almost one order of magnitude. Initial in-beam tests are scheduled at the RIBF in the near future NebulaPlus_RAPR . Physics opportunities such as the investigation of the two-neutron decay of 21B, the search for 21B, and the measurement of a three-neutron decay of 20B, are expected to be accessible with the NEBULA Plus detector.
@@
In contrast, a detector focused on more precise measurement by improving the achievable resolution in measurements of neutron momentum is also being constructed. The High-resolution Detector Array for Multi-neutron Events (HIME) detector was originally developed by the group of The Tokyo Institute of Technology (now Institute of Science Tokyo), and further collaborated with the Technical University of Darmstadt. The optimized energy resolution of HIME is achieved through the use of smaller scintillator bars with dimensions of 4 $\times$ 2 $\times$ 100 cm3, which improves position resolution, and new electronics for a higher time precision. It consists of layers of scintillator bars arranged in both the $x$ and $y$ directions alternatively Knoesel_thesis . This arrangement will not only improve its resolution, but will also allow discriminating the hits due to recoiling protons, allowing the selection of valid events and reducing cross-talk Hime_report . Currently, a prototype of HIME has been built, and more layers are in construction at TU Darmstadt. The improved energy resolution will enable precise measurements, such as the neutron-neutron scattering length, which characterizes the two-neutron interactions at low energies. Besides that, it is considered to be used for the detection of the recoil neutrons after the $(p,pn)$ reactions in combination with the proton trackers as described in the previous sections.
5 Conclusion
In conclusion, detector research and development (R&D) plays a crucial role in acquiring new physical data on quasi-free scattering using unstable nuclei. This review highlights several detectors that are either planned or under construction. One of the key projects is the STRASSE detection system, which is optimized for high-energy proton reactions and is under construction at the RIBF facility. It consists of a compact silicon tracker and a thick liquid hydrogen target, enabling high-resolution measurements of missing mass in proton-induced reactions. The system is designed to work in conjunction with the CATANA array, enhancing the detection of gamma rays and recoil protons with excellent energy resolution.
Another key project is the TOGAXSI telescope, developed as part of the ONOKORO project. It is designed to meet strict requirements for momentum transfer and angular coverage. With its advanced silicon trackers and GAGG(Ce) calorimeters, TOGAXSI is expected to significantly advance cluster research through high-precision missing mass spectroscopy.
In addition, in-beam $\gamma$-ray spectroscopy has proven essential for studying nuclear structure and excited states. The RIBF facility uses the DALI2 and its upgraded DALI2+ arrays to make groundbreaking discoveries, such as identifying the magic number at $N=34$ and exploring the ”Island of Inversion.” To fully benefit from upcoming RIBF upgrades, the next-generation HYPATIA array has been designed. It combines high-efficiency CeBr3 and HR-GAGG detectors in a modular, versatile configuration. This will provide better resolution, greater efficiency, and reduced background, ensuring continued progress in exploring exotic nuclei and their properties far from stability. Additionally, the future availability of a 238U primary beam at 345 MeV/nucleon with an intensity of 2000 pnA will enable the production of very exotic isotopes towards the drip lines. With its superior efficiency, time resolution, and peak-to-total ratio, the HYPATIA array will overcome previous limitations and make possible the first spectroscopy of 60Ca, along with other critical measurements for nuclear structure studies.
Furthermore, the state-of-the-art $\gamma$-ray tracking technique using Ge detectors has achieved significant breakthroughs in energy resolution. This is demonstrated by large-scale arrays such as GRETINA and AGATA. The GT-5 project, launched in 2023, aims to push the boundaries of in-beam $\gamma$-ray spectroscopy by integrating advanced tracking technology. It will provide systematic data to optimize optical models and enhance reaction analysis through collaborations with programs such as MESA and S3CAN.
Invariant mass spectroscopy is also proving to be an invaluable tool for understanding neutron-rich isotopes. The SAMURAI spectrometer and NEBULA detector are enabling the study of phenomena such as halo structures and unbound states. The development of NEBULA Plus and HIME will further improve neutron detection efficiency and energy resolution. This will open up new opportunities for precise measurements, such as the neutron-neutron scattering length, and advance our understanding of nuclear structure and neutron star properties.
@@ Acknowledgment
The authors acknowledge the contributions of Tomohiro Uesaka, Nigel Orr, Yasuhiro Togano, Marco Knösel, Valerii Panin, Ryotaro Tsuji, Rin Yokoyama, Pieter Doornenbal, Marina Petri, and Daisuke Suzuki. J.T. also acknowledges the ONOKORO collaboration.

### v3

@@ 3 Gamma-ray detection for gamma-ray spectroscopy
At the RIBF, secondary beams with incident energies of 100 to 250 MeV/u/u are used to populate excited states of nuclei far from stability. While such energies allow the use of thicker targets, and therefore increase the luminosity, they also present different experimental challenges. Of particular importance are the large Doppler shift of the detected $\gamma$-rays, and the background generated by either atomic or nuclear processes.
@@
where $\beta$ denotes the velocity of the beam and $\theta_{\gamma}$ is the angle between the particle direction and the $\gamma$-ray emission in the laboratory frame of reference. The energy resolution that can be achieved is therefore limited by the uncertainties in the velocity, $\Delta\beta$, the emission angle of the $\gamma$-ray, $\Delta\theta_{\gamma}$, and the intrinsic resolution of the detection material, $\Delta E_{\gamma}$. In general, the use of thicker secondary targets leads to a higher energy loss and therefore a larger contribution to $\Delta\beta$. The use of MINOS Obertelli_EPJA_2014 overcomes this problem by tracking the protons ejected from the target after knockout reactions and reconstructing the interaction position. In this way thicker targets can be used without affecting the achievable energy resolution as shown in Fig. 8. While the intrinsic resolution, $\Delta E_{\gamma}$, is a property of the detector material, the angular resolution, $\Delta\theta_{\gamma}$, can be improved either by reducing the size of the detectors or by implementing a position-sensitive readout with a sub-centimeter resolution for the first interaction of the $\gamma$-rays.
Figure 8: Total energy resolution (FWHM) for different $\gamma$-ray spectrometers. For a thick target, a large $\Delta\beta$ is obtained. This uncertainty is reduced by using MINOS. The use of smaller detectors with better intrinsic resolution, as the HYPATIA array, further improve the achievable energy resolution. The calculations were performed for $\beta=0.6$ and a 1 MeV $\gamma$-ray energy assuming a square-root dependence of the intrinsic energy resolution.
@@ 3.1 Towards a new scintillation detector array: The HYPATIA project
With this requirement in mind, the HYPATIA (HYbrid Photon detector Array To Investigate Atomic nuclei) array has been designed hypatia_web . This array, depicted in Fig. 9, consists of 624 Cerium Bromide (CeBr3) detectors of $30\times 30\times 80$ mm2 forming a barrel around the target and 384 Cerium-doped Gadolinium Aluminum Gallium Garnet (HR-GAGG) detectors of $25\times 25\times 100$ mm2 placed as a forward wall. The CeBr3 detectors were chosen due to their improved energy and time resolution compared to NaI(Tl), their reasonable cost, and the lack of intrinsic radiation. On the other hand, the HR-GAGG crystals are non-hygroscopic and have a high effective atomic number, therefore offering the possibility to minimize the dead space and increase the efficiency and peak-to-total ratio at forward angles. Thanks to the cuboidal shape of all the detectors, it is possible to arrange them in different configurations, including surrounding the MINOS or STRASSE targets, as well as to place the array at the different spectrometers of the RIBF. Furthermore, the selected shapes and sizes of the detectors are compatible with the mechanical structure of the DALI2+ array, therefore it will be possible to gradually replace the detectors without interrupting the physics program.
@@
Figure 10: a) Comparison between the performance of DALI2+ and HYPATIA for the measurement of 58Ca. b) Expected Doppler-corrected $\gamma$ spectrum of 60Ca for a 150-hours measurement with HYPATIA.
The Ca isotopic chain is particularly interesting for nuclear structure studies. While the proton shell closure at $Z=20$ seems to be maintained, new neutron shell closures at $N=32$ and $N=34$ have been evidenced by spectroscopy Huck_PhysRevC.31.2226 ; Steppenbeck_Nature_2013 , mass measurements Wienholtz_Nature2013 ; Michimasa_PhysRevLett.121.022506 , and knockout reaction cross sections Gade_PhysRevC.74.021302 ; Chen_PhysRevLett.123.142501 . The possibility of a further magic number at the $N=40$ harmonic-oscillator shell closure has been widely studied Gade_physics3040077 ; Brown_physics4020035 . Although measurements on 58Ca Chen_PLB_2023138025 and 62Ti Cortes_PLB_2020135071 , the closest neighbors of 60Ca, show no indication of a shell closure at $N=40$, only the first spectroscopy of 60Ca can provide definitive answers. This measurement will be possible with the HYPATIA array, as shown in Fig. 10b), where the simulated $\gamma$-ray spectrum obtained in a 150 h measurement is shown. As a reference, Fig. 10a) shows a comparison between the experimental spectrum obtained for 58Ca measured with DALI2+ Chen_PLB_2023138025 , and the expected performance of HYPATIA under the same conditions.
@@ 3.2 Future projects with Ge detectors
The $\gamma$-ray tracking technique of Ge detectors is the state-of-the-art method to achieve the best energy resolution for $\gamma$-rays. This technique has been implemented in the latest large-scale Ge arrays GRETINA PASCHALIS201344 and AGATA AKKOYUN201226 . During 2020 and 2021, the RIBF hosted the HiCARI array Wimmer2021 to conduct the first studies with tracking-type germanium detectors, which yielded successful outcomes. HiCARI was organized as a one-time campaign of experiments at the RIBF by gathering Ge detectors from other facilities. It is desirable to make continued efforts to push in-beam $\gamma$-ray spectroscopy based on tracking technology forward. For this reason, the Gamma-ray Tracking in R5 (GT-5) project was launched in 2023 at the RIKEN Nishina Center as a common and long-term platform for tracking-type Ge detectors at the RIBF to drive physics experiments and instrumentation for tracking technologies. The GT-5 has its own scientific and development programs related to in-beam $\gamma$-ray spectroscopy using RIBs. On a hardware basis, the GT-5 presently encompasses two in-house tracking detectors. One is the Compton camera telescope GREI motomura_IEEE_2007 , and the other is a position-sensitive strip germanium telescope SGT Suzuki_RAPR , both based on planar crystals with stripped segments. Furthermore, a GRETINA quad-type module was procured in 2023. In addition to the detectors, a maintenance facility for Ge detectors, electronics, and infrastructure is arranged. A unique aspect of the GT-5 is that it is not only involved in the scientific activities of the RIBF but also in the framework of the TRIP (Transformative Research Innovation Platform of RIKEN platforms) project, a RIKEN-wide initiative starting in 2023. One of the main goals of the TRIP use case is to establish an optical model for radioactive isotopes that are needed for reaction analyses and predictions but have yet to be studied based on systematic experimental data. To provide systematic data to optimize optical model potential parameters, the GT-5 is teamed with the elastic scattering measurement program MESA (Measurement of Elastic Scattering Anytime anywhere on any beam) and the reaction cross-section measurement program S3CAN (Symbiotic Systematic and Simultaneous Cross-section measurements for All over the Nuclear chart). As the first opportunities for data collection, experiments with stable beams and with RI beams have been conducted in 2024.
@@ 4 Neutron detection for invariant-mass spectroscopy
The upgrade plans of the RIBF will enable reaching of further neutron-rich isotopes, which are expected to have various weakly bound multi-neutron systems. To extend the sensitivity of multi-neutron reconstructions, a project to extend the detector volume for better detection efficiency, NEBULA Plus, has recently been completed through a collaborative effort between Japan and LPC Caen in France, as part of the French-funded EXPAND project expand . The NEBULA Plus array consisting of 90 new modules has been integrated with the existing array, effectively doubling the number of detection layers. This expansion increases the one-neutron detection efficiency by nearly a factor of 2 and improves the four-neutron detection efficiency by almost one order of magnitude. Initial in-beam tests were carried out at the RIBF in 2024 NebulaPlus_RAPR . Physics opportunities such as the investigation of the two-neutron decay of 21B, the search for 21B, and the measurement of a three-neutron decay of 20B, are expected to be accessible with the NEBULA Plus detector.
@@
In contrast, a detector focused on more precise measurement by improving the achievable resolution in measurements of neutron momentum is also being constructed. The High-resolution Detector Array for Multi-neutron Events (HIME) detector was originally developed by the group of The Tokyo Institute of Technology (now Institute of Science Tokyo), and further developed in collaboration with the Technical University of Darmstadt. The optimized energy resolution of HIME is achieved through the use of smaller scintillator bars with dimensions of 4 $\times$ 2 $\times$ 100 cm3, which improves position resolution, and new electronics for a higher time precision. It consists of layers of scintillator bars arranged alternating in $x$ and $y$ directions Knoesel_thesis . This arrangement will not only improve its resolution, but will also allow discriminating the hits due to recoiling protons, allowing the selection of valid events and reducing cross-talk Hime_report . Currently, a prototype of HIME has been built, and more layers are under construction at TU Darmstadt. The improved energy resolution will enable precise measurements, such as the neutron-neutron scattering length, which characterizes the two-neutron interactions at low energies. Besides that, it is planned to be used for the detection of the recoil neutrons after the $(p,pn)$ reactions in combination with the proton trackers as described in the previous sections.
5 Summary
Thanks to advances on the production of RIBs, isotopes that were previously inaccessible have been studied, and many interesting phenomena have been discovered. An integral part of research on the structure of exotic isotopes is the use of appropriate devices to detect the $\gamma$-rays, charged particles, and neutrons produced in the reactions.
A successful campaign of QFS experiments was performed at the RIBF using the MINOS LH2 target. The detection devices used for these experiments provided a plethora of results, and highlighted the limitations and possible improvements of the different detection systems. This experience, together with the future upgrade of the RIBF have promoted the development of new and improved detectors.
Based on the good performance of MINOS, the STRASSE Si tracker is under development. In addition to an improved resolution on the vertex reconstruction for knockout reactions, this system will provide the possibility to perform missing-mass measurement by the coupling with CATANA PLUS. A similar development, TOGAXI, will enable dedicated measurements of cluster knockout reaction. Thanks to the use of a calorimeter, missing mass-measurements will also be possible. Invariant mass measurements will benefit from the development of the neutron detectors NEBULA Plus and HIME, optimized for an increased efficiency and resolution, respectively. Furthermore, the study of bound excited states will profit from the developments on $\gamma$-ray detections arrays, such as HYPATIA, which will provide a better energy and time resolutions, as well as an improved peak-to-total ratio. Applications requiring optimized energy resolution can be addressed within the framework of the GT5 project.
In the future, the combination of QFS experiments with these new detection devices will help us to broaden our understanding of the structure of exotic isotopes. These detectors will also provide benchmarks for future developments in combination with the upgraded intesities of the RIBF.
@@ Acknowledgment
The authors acknowledge the contributions of Pieter Doornenbal, Marco Knösel, Alexandre Obertelli, Nigel Orr, Valerii Panin, Marina Petri, Daisuke Suzuki, Yasuhiro Togano, Ryotaro Tsuji, Tomohiro Uesaka, and Rin Yokoyama. J.T. acknowledges the ONOKORO collaboration and the support from the Grants-in-Aid for Scientific Research of the Japan Society for the Promotion of Science (JSPS) KAKENHI, Grant No. JP21H04975. M.L.C acknowledges the support of the RIKEN Special Postdoctoral Researcher Program. H.N.L thanks the STRASSE-CATANA collaboration, and acknowledges the support from the National Key R&D Program of China (Grant No. 2023YFA1606403, No. 2024YFE0102800), the National Natural Science Foundation of China (Grant No. 12375111), and the Fundamental Research Funds for the Central Universities of China. R.T acknowledges the support from the UK STFC and the Royal Society.